OpenStudy (anonymous):

5. Solve for x: 2x2 - 4x - 14 = 0 (completing the square)

OpenStudy (anonymous):

1-2squ 2 and 1+2squ2

OpenStudy (anonymous):

So far I have 2x^2 - 4 = 14 That can be factored so I get 2 (x^2 - 2x) = 14 Now i make my perfect square by dividing -2 by 2 and squaring that answer. Now I get 2 ( x^2 - 2x + 1) = 14 Then I factor it and get 2 (x - 1)^2 = 14 Now what do I do next?

OpenStudy (anonymous):

Then do I figure x - 1 = 7 to solve for x?

OpenStudy (anonymous):

wait... there's a problem... here: 2( x^2 - 2x + 1) = 14

OpenStudy (anonymous):

When I try to figure it out I keep getting x = 1, then I check it and its wrong. :(

OpenStudy (anonymous):

since you added 1 to the left side to make that perfect square trinomial, you actually added 2 to the left side. you must also add 2 to the right side.

OpenStudy (anonymous):

Divide each term by 2 to get: x² - 2x - 7 = 0 Complete the squares.. x² - 2x + c = 7 + c x² - 2x + 1 = 7 + 1 (x - 1)² = 8 Square root property: √(x - 1)² = ±√8 x - 1 = ±2√2 Hence, by adding both sides by 1, we obtain x = 1 ± 2√2

OpenStudy (anonymous):

but to make things simpler, i'd start out first by dividing the whole equation by 2 so i'd be working with \(\large x^2-2x-7=0 \) instead...

OpenStudy (anonymous):

^^^ that's it.... @jessica52595 did it for you.

OpenStudy (anonymous):

Thank you! @ByteMe and @jessica52595

OpenStudy (anonymous):

yw...:)